OpenStudy (anonymous):

Solve by the substitution method 3m+n=1 m-5n=11 Type the answer in an ordered pair.

OpenStudy (saifoo.khan):

m = 1 y= -2

OpenStudy (anonymous):

saifoo so the ordered pair looks like this? (1,-2)

OpenStudy (saifoo.khan):

Yes!

OpenStudy (anonymous):

grrrr...

OpenStudy (anonymous):

whoa totally different answers!

OpenStudy (anonymous):

uh oh...what is the grrr for?

OpenStudy (saifoo.khan):

chk yr answer, by substitutiong the values.,

OpenStudy (anonymous):

I get cranky. ;p You need to understand what was done and that it's correct in order to know what answer is right. So I hope you are actually working through these and if you have questions you ask.

OpenStudy (anonymous):

If 33+16n=1, then 16n=-32 so n=-2 Then when you sub in the other equation, the other value would also change. But Harkirat has the process correct.

OpenStudy (anonymous):

m - 5n = 11 => m = 11 + 5n In 3m + n = 1 substitute 11 + 5n for m to get 3(11 + 5n) + n = 1 33 + 15n + n = 1 16n = 1 - 33 n = -32/16 n = -2 now put n = -2 in m - 5n = 11 m - 5(-2) = 11 m + 10 = 11 m = 11 - 10 m = 1

OpenStudy (saifoo.khan):

np guyz, mistakes happen!

OpenStudy (anonymous):

sorry Rachel, i had made an error of sign at first, now the solution is correct n complete...

OpenStudy (anonymous):

I have thousands of problems some are actual homework and others are exercizes. I work through them on my own first and either I am not understanding because my homework says I am wrong or I am confirming my work.

OpenStudy (saifoo.khan):

np Harkirat.

OpenStudy (anonymous):

Thank you guys.

OpenStudy (anonymous):

thx saifoo n rachel n polpak :))